OpenStudy (anonymous):

Tell me if im right or wrong pleaseee???? Which is the completely factored form of x6 64y3? (Points : 4) (x^2- 4y)(x^4 + 4x^2y + 16y^2) (x^2 + 4y)(x^4- 4x^2y + 16y^2) (x^2 -4y)(x^4 -16y^2) (x^2 - 4y)^3 I think its B

OpenStudy (anonymous):

Which function is an even function? (Points : 4) f(x) = -x^2 f(x) = 6x^3 f(x) = -x- 2 f(x) = 2x^3 + 2 I think its B

OpenStudy (anonymous):

Which represents x2 - 7x - 60 in factored form? (Points : 4) (x -5)(x + 12) (x - 6)(x + 10) (x + 5)(x -12) (x - 5)(x - 12) I think the answer is C or D but im not sure???

OpenStudy (anonymous):

Which represents x^2 -12x + 36 in factored form? (Points : 4) (x + 6)(x -6) (x - 6)^2 (x + 6)^2 (x - 4)(x - 9) I think its B

OpenStudy (mathmath333):

for 1st question use formula \(\Large a^3-b^3=(a-b)(a^2+ab+b^2)\) \(\Large x^6-64y^3=(x^2)^3-(8y)^3\)

OpenStudy (mathmath333):

btw in ur question the sign is missing between Which is the completely factored form of x6 \( \Large (?)\) 64y3?"

OpenStudy (mathmath333):

a function is \(\large even\) function when \(\Large f(x)=f(-x)\) a function is \(\large odd\) function when \(\Large f(x)=-f(x)\)

OpenStudy (mathmath333):

\(\Large f(x) = 6x^3\) \(\Large f(-x) = 6(-x)^3=-6x^3\) \(\Large \text{so ur option B is not correct here}\) \

OpenStudy (mathmath333):

\(\Large f(x) = -x^2\) \(\Large f(-x) = -(-x)^2=-x^2\) so option \(\Large A\) is correct

OpenStudy (anonymous):

What about the other ones?

OpenStudy (mathmath333):

\(3.\) \(\Large x^2 - 7x- 60\) can be written as \(\Large x^2 - 12x+5x- 60\) now factor it

OpenStudy (anonymous):

x^2 -7x -60

OpenStudy (mathmath333):

last one is \(\Huge \checkmark\)

OpenStudy (anonymous):

I have 3 more can you check those @mathmath333

OpenStudy (anonymous):

Which is equal to (4x - 1)(3x + 2)? (Points : 4) 12x^2 - 5x + 2 12x^2 + 5x - 2 7x^2 + 5x - 2 12x^2 - 2 I think its A

OpenStudy (anonymous):

Which is equal to (2x^2 - 5x + 4) (4x^2 + 3x - 2)? (Points : 4) 2x^2- 8x + 6 6x^2 - 2x + 2 2x^2- 2x + 2 2x^4 - 8x^2 + 6 I think its C

OpenStudy (anonymous):

Which represents the polynomial in standard form? 12x^2 - 9x^3 + 5x^2 - 6x^3 + 3 + 2x (Points : 4) 17x^2 - 15x^3 + 2x + 3 4x^11 + 3 -3x^3 + 17x^2 + 2x + 3 -15x^3 + 17x^2 + 2x + 3 I think its D

OpenStudy (anonymous):

@mathmath333 could you please check these?

OpenStudy (anonymous):

@FireWolfSpirit can you check them ?

OpenStudy (mathmath333):

\(\Large (4x - 1)(3x + 2)\) its not option \(\Large A\) try again

OpenStudy (mathmath333):

i think the question is in correct here Which is equal to (2x^2 - 5x + 4) (4x^2 + 3x - 2)?

OpenStudy (mathmath333):

or the options

OpenStudy (mathmath333):

Which represents the polynomial in standard form? 12x^2 - 9x^3 + 5x^2 - 6x^3 + 3 + 2x option \(\Large D\) is correct here